\underline{1 USAGE}

Example for converting a L^{A}TEX document to

                 HTML

using L^{A}TEX2HTML and the rcsinfo-package

Dr. Jürgen Vollmer Viktoriastraße 15, D-76133 Karlruhe, Germany

                         Juergen.Vollmer@acm.org

February 5, 1999

1 Usage

Note: If using the L^{A}TEX2HTML tool, you must place \rcsInfo into the preamble too, i.e. before the \begin{document}. Using only L^{A}TEX, \rcsInfo must be placed after the preamble.

Note: If the fancy / fancyhdr option is given to the rcsinfo package, the date of the ADDRESS at the bottom of the HTML page is set to the RCS-date.

Note: If you would like to set your own ADDRESS text, you should define a procedure, which sets the perl-variable $main::ADDRESS in your local .latex2html initialization file. The routine may use the perl variables shown in the section 2. Assign a reference of this procdure to the perl-variable

$rcsinfo::SetAddressProc.

For example the follownig perl-code does the job:

sub my_address_field
{
 $ADDRESS = '<hr>This file was last updated on ' .
            $rcsinfo::Date  .
            '&nbsp; by ' .
            '<a href="mailto: Juergen.Vollmer@acm.org">' .
            'Juergen Vollmer ' .
    '&lt;Juergen.Vollmer@acm.org&gt;</a>';
}
$rcsinfo::SetAddressProc = \&my_address_field;

\underline{2 EXAMPLE}

2 Example

Here are the commands and their output:

L^{A}TEX-command perl variable Result for this file
\today            $today  February 5, 1999

\rcsInfoFile      $rcsinfo::File  rcsinfo.doc
\rcsInfoOwner     $rcsinfo::Owner  vollmer
\rcsInfoStatus    $rcsinfo::Status  Entwurf
\rcsInfoLocker    $rcsinfo::Locker  vollmer
\rcsInfoDate      $rcsinfo::Date  1999/02/05
\rcsInfoLongDate  $rcsinfo::LongDate February 5, 1999
\rcsInfoDay       $rcsinfo::Day  5
\rcsInfoMonth     $rcsinfo::Month  2
\rcsInfoYear      $rcsinfo::Year  1999